# Quantum Polymorphisms and Commutativity Gadgets

### v1

@@ 8.2. The minimal contextual clone
Consider a relational structure $\mathbb{A}:=(\{0,1\};R)$ that is preserved by majority, then it is well known that $R$ (of arity $r$) can be pp-defined in $\mathbb{A}_{2}:=(\{0,1\};\{\pi_{i,j}(R)\mid 1\leq i,j\leq r\})$ via the conjunction of its binary projections $\bigwedge_{1\leq i,j\leq r}\pi_{\{i,j\}}(R)(x_{i},x_{j})$. We show that $R$ is preserved by every quantum polymorphism $Q\colon\mathbb{A}_{2}^{n}\to\mathbb{A}_{2}$. Assume that $(\mathbf{a}_{1},\dots,\mathbf{a}_{r})$ is in the $n$th direct power of $R$ while in $(x_{1},\dots,x_{r})$, we have indices $i,j$ such that $(x_{i},x_{j})\notin\pi_{\{i,j\}}(R)$. Then
@@
where the second equality is by duplicating the $Q_{\mathbf{a}_{i},x_{i}}$ and $Q_{\mathbf{a}_{j},x_{j}}$ projectors and then using property $($$\mathbf{QH_{2}}$)) to move them to the front. Moreover, if $\mathbf{a},\mathbf{a}^{\prime}\in E(\operatorname{Gaif}(\mathbb{A}^{n}))$ then $\mathbf{a},\mathbf{a}^{\prime}\in E(\operatorname{Gaif}(\mathbb{A}_{2}^{n}))$ so, for any $x,x^{\prime}\in\{0,1\}$ $Q_{\mathbf{a},x},Q_{\mathbf{a}^{\prime},x^{\prime}}]=0$ by property $($$\mathbf{QH_{2}}$)). $Q$ is then an oracular quantum polymorphism of $\mathbb{A}$ and so $\operatorname{qPol}(\mathbb{A}_{2})\subseteq\operatorname{qPol}(\mathbb{A})$.
By assumption, no binary projection of a relation of $\mathbb{A}$ is full, thus, each relation $R$ in $\mathbb{A}_{2}$ can be defined as the (non-empty!) conjunction $\bigwedge_{(c,d)\not\in R}S_{cd}(x,y)$, where we use the shorthand $S_{01}(x,y)$ for the formula $S_{10}(y,x)$. We show that each relation $R$ of $\mathbb{A}_{2}$ is preserved by the quantum polymorphisms of $\mathbb{B}$.
Let $Q\colon\mathbb{B}^{n}\to_{q}\mathbb{B}$ be a quantum polymorphism, $(\mathbf{a},\mathbf{b})$ be in the $n$th direct power of $R$ and $(c,d)\notin R$. We then have $Q_{\mathbf{a},c}Q_{\mathbf{b},d}=0$ by $($$\mathbf{QH_{1}}$)) since $(c,d)\not\in S_{cd}$ and $(\mathbf{a},\mathbf{b})$ is in the relation $S_{cd}$ in the $n$th power of $\mathbb{B}$. Thus, $($$\mathbf{QH_{1}}$)) holds.
Finally, if $(\mathbf{a},\mathbf{b})$ is a pair in a relation $R$ in the $n$th power of $\mathbb{A}_{2}$, then it is necessarily a pair in one of the relations $S_{cd}$ in the $n$th power of $\mathbb{B}$. Thus, $Q_{\mathbf{a}}$ and $Q_{\mathbf{b}}$ commute, so that $($$\mathbf{QH_{2}}$)) holds. ∎
Now we show that this clone actually contains some contextual polymorphisms.
@@
For $n\leq 3$, every quantum polymorphism $Q\colon\mathbb{B}^{n}\to\mathbb{B}$ is non-contextual. For all $n\geq 4$, $\mathbb{B}$ has contextual quantum polymorphisms of arity $n$.
@@
We start by proving that every quantum polymorphism $Q\colon\mathbb{B}^{n}\to_{q}\mathbb{B}$ is non-contextual for $n\leq 3$. For this, it suffices by $($$\mathbf{QH_{2}}$)) to prove that for every pair $S,T$ of subsets of $\{1,\dots,n\}$, either $(S,T)$ or $(T,S)$ belongs to one of the relations of $\mathbb{B}^{n}$. If $(S,T)$ does not belong to the relation $S_{00}$ in $\mathbb{B}^{n}$, then there exists $i\in\{1,\dots,n\}$ such that $i\in\overline{S}\cap\overline{T}$. Similarly, for it to not belong to the relation $S_{11}$ in $\mathbb{B}^{n}$, there must exist $j\in\{1,\dots,n\}$ such that $j\in S\cap T$. If $n\leq 3$, there remains at one most other element $k$ in $\{1,\dots,n\}\setminus\{i,j\}$. If $k$ is not in $S\cap T$ nor in $\overline{S}\cap\overline{T}$, then it is the case that $(S,T)$ or $(T,S)$ belongs to the relation $S_{10}$ of $\mathbb{B}^{n}$, so we are done.
For the $n\geq 4$ case we define a concrete quantum polymorphism for $n=4$. We provide Figure 3 as a visual aid.
Figure 3. Simplified illustration of $\mathbb{B}^{4}$. Black lines indicate the poset $S_{10}$. The blue (resp. red) dotted lines show parts of $S_{00}$ (resp. $S_{11}$); this relation additionally relates $(0,0,0,0)$ (resp. $(1,1,1,1)$) to everything and relates two nodes when they are a bit-flips of each other (see coloring).
We define $Q\colon\mathbb{B}^{4}\to_{q}\mathbb{B}$ by listing the projectors. Let $T\subseteq[4]$. If $|T|\geq 3$, define $Q_{T}=\operatorname{id}$. Pick now arbitrary matrices $A,B,C$. Define $Q_{\{1,2\}}:=A$, $Q_{\{1,3\}}=B$, and $Q_{\{1,4\}}=C$. In all the other cases, define $Q_{T}=\operatorname{id}-Q_{[4]\setminus T}$. Note that if at least two of the matrices $A,B,C$ do not commute, then $Q$ is contextual.
It remains to prove that $Q$ preserves the relations of $\mathbb{B}^{4}$. Assume we have $S,T\subseteq[4]$ with $(s,t)$ in the fourth power of $S_{11}$ and $x,y\in\{0,1\}$ with $(x,y)\notin S_{11}$. If $S\neq\overline{T}$, we are in one of 2 cases; either $S$ or $T$ is the empty set or $|S|=|T|=1$. In both cases at least $Q_{S}$ or $Q_{T}$ equals 0 and we are done immediately. If $S=\overline{T}$, then $Q_{S}Q_{T}=Q_{S}Q_{\overline{S}}=Q_{S}(\operatorname{id}-Q_{S})=0$ and we are done. The case of $S_{00}$ is symmetric.
Now assume we have $S,T\subseteq[4]$ such that the corresponding pair of tuples $(s,t)$ in the fourth power of $S_{10}$. If both $|S|$ and $|T|$ do not equal $2$ we know $Q_{S},Q_{T},(\operatorname{id}-Q_{S}),(\operatorname{id}-Q_{T})\in\{0,\operatorname{id}\}$ and we immediately get that $Q_{S},{\operatorname{id}-Q_{T}}]=0$. The final case is when one of $|S|$ or $|T|$ equal $2$, here we get that either
@@
showing preservation of the relations. ∎
9. Acknowledgments

### v2

@@ 11.2 The minimal contextual minion
Consider a relational structure $\mathbb{A}:=(\{0,1\};R)$ that is preserved by majority, then it is well known that $R$ (of arity $r$) can be pp-defined in $\mathbb{A}_{2}:=(\{0,1\};\{\pi_{i,j}(R)\mid 1\leq i,j\leq r\})$ via the conjunction of its binary projections $\bigwedge_{1\leq i,j\leq r}\pi_{\{i,j\}}(R)(x_{i},x_{j})$. We show that $R$ is preserved by every quantum polymorphism $Q\colon\mathbb{A}_{2}^{n}\to_{\operatorname{q}}\mathbb{A}_{2}$. Assume that $(\mathbf{a}_{1},\dots,\mathbf{a}_{r})$ is in the $n$th direct power of $R$ while in $(x_{1},\dots,x_{r})$, we have indices $i,j$ such that $(x_{i},x_{j})\notin\pi_{\{i,j\}}(R)$. Then
@@
where the second equality is by duplicating the $Q_{\mathbf{a}_{i},x_{i}}$ and $Q_{\mathbf{a}_{j},x_{j}}$ projectors and then using property $($$\mathbf{QH_{2}}$)) to move them to the front. Moreover, if $(\mathbf{a},\mathbf{a}^{\prime})\in E(\operatorname{Gaif}(\mathbb{A}^{n}))$ then $(\mathbf{a},\mathbf{a}^{\prime})\in E(\operatorname{Gaif}(\mathbb{A}_{2}^{n}))$ so, for any $x,x^{\prime}\in\{0,1\}$ $Q_{\mathbf{a},x},Q_{\mathbf{a}^{\prime},x^{\prime}}]=0$ by property $($$\mathbf{QH_{2}}$)). $Q$ is then an oracular quantum polymorphism of $\mathbb{A}$ and so $\operatorname{qPol}(\mathbb{A}_{2})\subseteq\operatorname{qPol}(\mathbb{A})$.
By assumption, no binary projection of a relation of $\mathbb{A}$ is full, thus, each relation $R$ in $\mathbb{A}_{2}$ can be defined as the (non-empty) conjunction $\bigwedge_{(c,d)\not\in R}S_{cd}(x,y)$, where we use the shorthand $S_{01}(x,y)$ for the formula $S_{10}(y,x)$. We show that each relation $R$ of $\mathbb{A}_{2}$ is preserved by the quantum polymorphisms of $\operatorname{\mathbb{M}}$.
Let $Q\colon\operatorname{\mathbb{M}}^{n}\to_{\operatorname{q}}\operatorname{\mathbb{M}}$ be a quantum polymorphism, $(\mathbf{a},\mathbf{b})$ be in the $n$th direct power of $R$ and $(c,d)\notin R$. We then have $Q_{\mathbf{a},c}Q_{\mathbf{b},d}=0$ by $($$\mathbf{QH_{1}}$)) since $(c,d)\not\in S_{cd}$ and $(\mathbf{a},\mathbf{b})$ is in the relation $S_{cd}$ in the $n$th power of $\operatorname{\mathbb{M}}$. Thus, $($$\mathbf{QH_{1}}$)) holds.
Finally, if $(\mathbf{a},\mathbf{b})$ is a pair in a relation $R$ in the $n$th power of $\mathbb{A}_{2}$, then it is necessarily a pair in one of the relations $S_{cd}$ in the $n$th power of $\operatorname{\mathbb{M}}$. Thus, $Q_{\mathbf{a}}$ and $Q_{\mathbf{b}}$ commute, so that $($$\mathbf{QH_{2}}$)) holds. ∎
Now we show that the minion defined here, actually contains some contextual polymorphisms.
@@
For $n\leq 3$, every quantum polymorphism $Q\colon\operatorname{\mathbb{M}}^{n}\to\operatorname{\mathbb{M}}$ is non-contextual. For all $n\geq 4$, $\operatorname{\mathbb{M}}$ has contextual quantum polymorphisms of arity $n$.
@@
We start by proving that every quantum polymorphism $Q\colon\operatorname{\mathbb{M}}^{n}\to_{\operatorname{q}}\operatorname{\mathbb{M}}$ is non-contextual for $n\leq 3$. For this, it suffices by $($$\mathbf{QH_{2}}$)) to prove that for every pair $S,T$ of subsets of $\{1,\dots,n\}$, either $(S,T)$ or $(T,S)$ belongs to one of the relations of $\operatorname{\mathbb{M}}^{n}$. If $(S,T)$ does not belong to the relation $S_{00}$ in $\operatorname{\mathbb{M}}^{n}$, then there exists $i\in\{1,\dots,n\}$ such that $i\in\overline{S}\cap\overline{T}$. Similarly, for it to not belong to the relation $S_{11}$ in $\operatorname{\mathbb{M}}^{n}$, there must exist $j\in\{1,\dots,n\}$ such that $j\in S\cap T$. If $n\leq 3$, there remains at one most other element $k$ in $\{1,\dots,n\}\setminus\{i,j\}$. If $k$ is not in $S\cap T$ nor in $\overline{S}\cap\overline{T}$, then it is the case that $(S,T)$ or $(T,S)$ belongs to the relation $S_{10}$ of $\operatorname{\mathbb{M}}^{n}$, so we are done.
We now build a concrete class of contextual polymorphisms of arity $n=4$. We provide Figure 2 as a visual aid in this example. The figure is a simplified presentation of the relatedness of tuples in $\{0,1\}^{4}$ with the arrows indicating the poset $S_{10}$ and the colouring emphasizing that $S_{11}$ (and $S_{00}$) relate all opposite tuples. In particular, the only unrelated tuples are those $s,t$ such that $|S|=|T|=2$ and $S\neq\overline{T}$.
We define $Q\colon\operatorname{\mathbb{M}}^{4}\to_{\operatorname{q}}\operatorname{\mathbb{M}}$ by listing the projectors. Let $T\subseteq[4]$. If $|T|\geq 3$, define $Q_{T}=\operatorname{id}$. Pick now arbitrary matrices $A,B,C$. Define $Q_{\{1,2\}}:=A$, $Q_{\{1,3\}}=B$, and $Q_{\{1,4\}}=C$. In all the other cases, define $Q_{T}=\operatorname{id}-Q_{[4]\setminus T}$. Note that if at least two of the matrices $A,B,C$ do not commute, then $Q$ is contextual.
It remains to prove that $Q$ preserves the relations of $\operatorname{\mathbb{M}}^{4}$. Assume we have $S,T\subseteq[4]$ with $(s,t)$ in the fourth power of $S_{11}$ and consider $Q_{S}Q_{T}$ (since $(1,1)$ is the only Boolean pair not in $S_{11}$). $S$ and $T$ are related if, and only if, $S\cap T=\emptyset$; hence, we have two cases: Either $|S|=|T|=2$ and $S=\overline{T}$, or one of $S,T$ has cardinality less than $2$.
If $S\neq\overline{T}$, then $|S|\leq 1$ or $|T|\leq 1$ in this case at least $Q_{S}$ or $Q_{T}$ equals 0 and we are done immediately. If $S=\overline{T}$, then $Q_{S}Q_{T}=Q_{S}Q_{\smash{\overline{S}}}=Q_{S}(\operatorname{id}-Q_{S})=0$ and we are done. The case of $S_{00}$ is symmetric.
Now assume we have $S,T\subseteq[4]$ such that the corresponding pair of tuples $(s,t)$ in the fourth power of $S_{10}$ ($S\subseteq T$). If both $|S|$ and $|T|$ do not equal $2$, we know that all projectors $Q_{S}$, $Q_{T}$, $(\operatorname{id}-Q_{S})$, and $(\operatorname{id}-Q_{T})$ are either 0 or $\operatorname{id}$; thus, we immediately find that $Q_{S},{\operatorname{id}-Q_{T}}]=0$. The final case is when at least one of $|S|$ or $|T|$ equals $2$. In this case, we deduce that either
@@
when $|S|=2$ (and $S\neq T$). Thus all relations are preserved, and $Q\colon\operatorname{\mathbb{M}}^{4}\to_{\operatorname{q}}\operatorname{\mathbb{M}}$ is a quantum polymorphism.
Note that one can easily extend $Q$ to a contextual polymorphism of arbitrary arity $\geq 4$ by adding dummy variables. For example, define $\tau\colon[4]\to[n]$ by $\tau(i):=i$ for $i\in[4]$. By Proposition 26,
$Q_{/\tau}=Q\bullet\bigotimes_{j\in[4]}\pi_{\tau(j)}^{n}\colon M^{n}\to_{\operatorname{q}}M$
is in $\operatorname{qPol}(\operatorname{\mathbb{M}})$. We just need to check that $Q_{/\tau}$ is non-contextual, but this is immediate from the definition of $Q$, since e.g. $(Q_{/\tau})_{\{1,2\}}=Q_{\{1,2\}}=A$, $(Q_{/\tau})_{\{1,3\}}=Q_{\{1,3\}}=B$, and $A,B]\neq 0$. ∎
We can now conclude by proving the main result on this section, giving a full classification of the existence of commutativity gadgets for Boolean structures. As remarked in Section 2.6, the $(ii)\Rightarrow(i)$ implication was first proved in [33], which described explicit commutativity gadgets for $\mathbb{A}$ having no majority and for any $\mathbb{A}$ that is not TVF.
Proof of Theorem 11.
  • $(ii)\Rightarrow(i)$
    If $\mathbb{A}$ has no majority polymorphism, then it admits a commutativity gadget by Theorem 114 and Theorem 6. Otherwise, the result follows from Proposition 61 and Theorem 6.
  • $(i)\Rightarrow(ii)$
    By contrapositive, if $\mathbb{A}$ admits a majority polymorphism and is TVF, then it holds that $\operatorname{qPol}(\operatorname{\mathbb{M}})\subseteq\operatorname{qPol}(\mathbb{A})$ by Theorem 115. Thus $\mathbb{A}$ has at least one contextual polymorphism by Theorem 116. So, it has no commutativity gadget.∎
12 Acknowledgments